### v1

@@ 6 Conclusions
IRKGL methods are symplectic and, in non-stiff regimes, can be efficiently implemented using fixed-point iterations. These methods combine symplecticity, time-reversibility, and high-order of accuracy, and they are broadly applicable to general canonical Hamiltonian systems [18]. While optimized high-order explicit symplectic integrators are effective when the Hamiltonian can be decomposed into solvable components, they may be less efficient than IRKGL methods when high-precision integrations are required. The SIMD-based implementation of IRKGL16 significantly enhances performance, often outperforming specialized explicit symplectic methods even at moderate precision levels.
For non-stiff problems with Hamiltonians of the specific form (17), optimized high-order explicit symplectic integrators remain highly effective due to their simplicity and computational efficiency. Although our sequential implementation of IRKGL16 is less efficient than the best available explicit symplectic methods for standard double-precision computations, the SIMD-enhanced version can surpass these methods in performance for high-accuracy integrations, even when constrained to double-precision arithmetic.
@@ Acknowledgments
Appendix A Efficient implementation of the three-part splitting method for the Schwarzschild black hole example
The evaluation of the flows $\varphi_{t}^{A}$ and $\varphi_{t}^{B}$ becomes more efficient when performed in the variables $(x,y,p_{x},p_{y})$, which are related to the original coordinates by the canonical transformation
$x$ $=r\cos\theta,\quad y=r\sin\theta,$
$p_{x}$ $=\cos\theta\,p_{r}-\sin\theta\,\frac{p_{\theta}}{r},\quad p_{y}=\sin\theta\,p_{r}+\cos\theta\,\frac{p_{\theta}}{r}.$
In these new variables, the Hamiltonians $H_{A}$ and $H_{B}$ become
$\bar{H}_{A}(x,y)$ $=\frac{1}{2y^{2}}\left(L-\frac{\beta}{2}y^{2}\right)^{2}-\frac{1}{2}\left(1-\frac{2}{\sqrt{x^{2}+y^{2}}}\right)^{-1}E^{2},$
$\bar{H}_{B}(p_{x},p_{y})$ $=\frac{1}{2}(p_{x}^{2}+p_{y}^{2}).$
For the $H_{C}$ component, the equations of motion are simplified by introducing the variables $(\nu,\mu)$, defined through the non-canonical transformation
$\nu=r\,p_{r},\quad\mu=\frac{p_{r}^{2}}{r}.$
The corresponding equations are
$\frac{d\nu}{dt}=-3\mu,\quad\frac{d\mu}{dt}=0.$
The flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$—and hence the integrators (22) and (23)—can be efficiently evaluated in the variables $(x,y,p_{x},p_{y})$ as follows:
  • 1)
    Flow of $H_{A}$:
    $\varphi_{t}^{A}:(x,y,p_{x},p_{y})\mapsto(x,y,p_{x}^{*},p_{y}^{*}),$
    where
    $p_{x}^{*}$ $=p_{x}-t\,\frac{\partial\bar{H}_{A}}{\partial x}(x,y),$
    $p_{y}^{*}$ $=p_{y}-t\,\frac{\partial\bar{H}_{A}}{\partial y}(x,y).$
  • 2)
    Flow of $H_{B}$:
    $\varphi_{t}^{B}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x},p_{y}),$
    where
    $x^{*}$ $=x+t\,p_{x},\quad y^{*}=y+t\,p_{y}.$
  • 3)
    Flow of $H_{C}$:
    $\varphi_{t}^{C}:(x,y,p_{x},p_{y})\mapsto(x^{*},y^{*},p_{x}^{*},p_{y}^{*}),$
    where
    $r$ $=\sqrt{x^{2}+y^{2}},\quad c=\frac{x}{r},\quad s=\frac{y}{r},$
    $p_{\theta}$ $=-y\,p_{x}+x\,p_{y},\quad\nu=x\,p_{x}+y\,p_{y},\quad\mu=\frac{\nu^{2}}{r^{3}},$
    $\nu^{*}$ $=\nu-3t\,\mu,\quad p_{r}^{*}=(\mu\,\nu^{*})^{1/3},\quad r^{*}=\frac{\nu^{*}}{p_{r}^{*}},$
    $x^{*}$ $=r^{*}\,c,\quad y^{*}=r^{*}\,s,$
    $p_{x}^{*}$ $=c\,p_{r}^{*}-s\,\frac{p_{\theta}}{r^{*}},\quad p_{y}^{*}=s\,p_{r}^{*}+c\,\frac{p_{\theta}}{r^{*}}.$
The partial derivatives of $\bar{H}_{A}(x,y)$ can be computed efficiently using the intermediate variables:
$z$ $=y^{2},\quad r=\sqrt{x^{2}+z},\quad w=L-\frac{\beta}{2}z,\quad u=1-\frac{2}{r},$
$W$ $=-\frac{w}{z},\quad U=-\frac{E^{2}}{2u^{2}},\quad R=\frac{U}{r^{3}},\quad Z=R+\frac{(W-\beta)W}{2},$
$\frac{\partial\bar{H}_{A}}{\partial x}$ $=-2xR,\quad\frac{\partial\bar{H}_{A}}{\partial y}=-2yZ.$
Unlike the analytical expressions provided in [15] for the flows $\varphi_{t}^{A}$, $\varphi_{t}^{B}$, and $\varphi_{t}^{C}$, our implementation avoids the use of trigonometric and inverse trigonometric functions entirely, resulting in improved computational efficiency.

### v2

@@ 6 Conclusions
High-order explicit symplectic integrators are effective when the Hamiltonian can be decomposed into solvable components, but IRKGL-based integrators offer a powerful alternative: they are symplectic, time-reversible, and high-order accurate, can be efficiently implemented using fixed-point iterations in non-stiff regimes, and are broadly applicable to general canonical Hamiltonian systems [21, 13] in first-order ODE form, regardless of whether the Hamiltonian can be split into solvable parts.
In Section 2, we presented a general stage-wise vectorized formulation of IRKGL schemes for both first-order and second-order Hamiltonian systems. Beyond reducing computational overhead, this formulation leaves the algorithm naturally ready for stage-wise SIMD vectorization, enabling parallel evaluation of the right-hand side function across all stages whenever the function is compatible with SIMD.jl. For second-order systems, we introduced a reformulation of the corresponding Runge–Kutta–Nyström method that ensures exact symplecticity at the level of double-precision floating-point arithmetic, extending the approach of [1] originally developed for first-order systems to this setting.
The numerical experiments of Section 5 confirm the effectiveness of the proposed approach. For Hamiltonian problems admitting a decomposition into a sum of exactly solvable components, IRKGL16 outperforms optimized high-order explicit symplectic integrators in the high-precision regime across all test problems considered. When the right-hand side function is compatible with SIMD.jl, this advantage is most pronounced; when it is not, the stage-wise vectorization of all remaining computations still greatly reduces overhead, and IRKGL16 may still surpass explicit methods at high-precision levels.
Overall, our results show that IRKGL integrators, optimized with the proposed stage-wise SIMD vectorization technique, provide a powerful and general alternative to explicit symplectic methods, achieving high accuracy efficiently for non-stiff Hamiltonian systems.